Is the population in Hardy Weinberg Equilibrium? Show your work.


Generation 1: p=0.6, q=0.4; population mean fitness =0.9 

Generation 2: p'=0.54, q'=0.46 


The genotype frequency in the first generation is: 

AA: 0.3 

Aa: 0.5 

aa: 0.2 

Is the population in Hardy Weinberg Equilibrium? Show your work.  
Answer #1

Given

Allele frequency of A = 0.6

Allele frequency of a = 0.4

Genotype frequency

AA = 0.3

Aa = 0.5

aa = 0.2

Accordint to hardy weinberg population

Genotye frequency of AA = p2 = 0.6*0.6 = 0.36

Genotye frequency of aa = q2 = 0.4*0.4 = 0.16

Genotye frequency of Aa = 2pq = 2*0.6*0.4= 0.48

This is almost similar to given genotype frequency.

So the population is in hardy Weinberg equilibrium.